# Johannes Matzat        Aiko Schmeißer

### v4

@@ 4 Empirical Strategy
Our objective is to estimate the causal effect of unionization on employees’ campaign contributions. A simple comparison of employees in unionized and non-unionized workplaces will fail to account for differences between these groups along a number of dimensions. These arise because the decision to unionize is likely endogenous and correlated with many characteristics, among them potentially political behavior. Figure 1 depicts average campaign contribution amounts across winning and losing union elections before and after the election. Due to their shared interest in a union election at the same time, these establishments are expected to be more similar than a random sample of unionized and non-unionized establishments. [^23] Dinlersoz et al. (2017) examine selection into union elections and find that elections are more likely to be held at younger, larger, more productive, and higher-paying establishments. Our strategy avoids such selection by comparing only establishments that hold union elections. Pre-existing ideological differences are nevertheless visible: Workplaces that vote for unionization donate more to Democratic candidates and less to Republican candidates even before the union election.
@@
To account for pre-existing differences, we implement a difference-in-differences approach and compare campaign contribution patterns before and after the union election in establishments where the union won versus where it lost. We complement the DiD design with methods from the RDD literature to probe the validity of the underlying parallel trends assumption. In particular, we exploit the fact that we observe the pro-union vote share, which discontinuously determines unionization at the 50% threshold. We use the vote share to estimate placebo tests for differential trends by vote shares among losing union elections as well as to examine the robustness of our DiD estimates when restricting the sample to establishments with increasingly close election results. [^24] Many papers on the effects of unionization follow an RDD by comparing establishments in which the union barely won versus where it barely lost (e.g., Campello et al., 2018; DiNardo and Lee, 2004; Ghaly et al., 2021; Lee and Mas, 2012; Sojourner et al., 2015; Sojourner and Yang, 2022). This approach is complicated by the fact that unions and employers can influence election outcomes even after the election, through challenging the validity of individual ballots or filing charges of unfair labor practices. Frandsen (2021) and Knepper (2020) provide evidence for discontinuities at the 50% threshold in the vote share distribution and in pre-election establishment characteristics. Appendix Figure A.2 verifies that, also in our matched sample of elections, there is a significant discontinuity in the vote share density at the 50% cutoff, indicating a manipulation of closely contested elections. Finally, to cross-validate the causal interpretation of our results, we also develop a novel identification strategy that combines the DiD with an IV approach. For this, we exploit variation in unionization resulting from exogenous shocks to the salience of safety at work that are triggered by unexpected fatal workplace accidents shortly before the union election.
@@
where $y_{jk}$ denotes a political outcome for employees in establishment $j$ and relative event time $k$. We observe each establishment from three cycles before to three cycles after the union election, i.e., $k=\{-3,-2,...,3\}$, where $k=0$ refers to the cycle in which the union election takes place. Our effect of interest is captured by $\delta_{\textrm{DiD}}$. It is the coefficient of an interaction term between a post-treatment dummy and a dummy indicating whether the election was won by the union, i.e., whether the pro-union vote share, $V_{j}$, is above 50%. $\alpha_{j}$ denotes establishment fixed effects that capture all time-invariant differences between winning and losing establishments. Further, we introduce event-time $\times$ cohort fixed effects $\beta_{kc_{j}}$, where cohort $c_{j}$ refers to the political election cycle in which the union election was held, i.e., $c_{j}=\{1985/86,1987/88,...,2009/10\}$. Importantly, with these fixed effects our identifying variation only comes from comparing changes across winning and losing elections within the same cohort. Thereby, it avoids “forbidden comparisons” between late and early-treated establishments that may lead to negative weights when averaging potentially heterogeneous, cohort-specific treatment effects in staggered DiD settings such as ours (de Chaisemartin and D’Haultfœuille, 2020; Goodman-Bacon, 2021; Sun and Abraham, 2021). Our DiD model is equivalent to the stacking approach first implemented by Cengiz et al. (2019). This approach first creates cohort-specific datasets of treated units and an appropriate set of control units that are never or not yet treated. Then, one stacks the cohort-specific datasets by time relative to treatment start in order to estimate an average treatment effect across all cohorts. By stacking and aligning cohorts in relative time, this strategy mimics a setting where all treatments occur contemporaneously and thus avoids using already-treated units in the comparison group. Note that in our case the selection of appropriate control units for the stacking is facilitated by the possibility that we can naturally compare treated establishments to untreated establishments that have a lost election in the same cycle. Finally, we cluster standard errors at the level of treatment, the establishment.
Model (1) pools all periods after treatment, which yields the maximum power when estimating average treatment effects. To examine how treatment effects vary by event time, we also estimate the following stacked event-study model:
@@
where the $\delta_{s}$ coefficients capture dynamic treatment effects relative to the cycle before the union election was held (the interaction with $k=-1$ is omitted).
Parallel trends assumption. Our identifying assumption is that campaign contributions for winning establishments would have evolved in parallel to contributions in losing establishments had the union not won the election:
[…]
@@
We run different tests to examine the validity of this assumption. First, we analyze whether outcomes developed in parallel before the election. Figure 1 provides first visual evidence that pre-election changes in contribution amounts to Republican and Democratic candidates are very similar across winning and losing elections. [^25] Note that the strong upward trends in both winning and losing elections are explained by the increasing importance of campaign contributions in more recent election campaigns. The pre-election $\delta_{s}$ coefficients estimated in the event study model will provide a formal test of pre-trends.
Second, even in the absence of significant pre-trends, there may still be unobserved shocks that drive union voting results at the time of the election and that may be related to changes in contribution patterns. To test whether such shocks likely violate our identifying assumption, we follow the approach of Wang and Young (2023) and analyze whether changes in outcomes are different among losing elections with different vote shares. If unobserved shocks were driving voting results that led to union victory or loss, we would also expect them to affect outcomes in losing elections with different union vote shares. [^26] Wang and Young (2023) formulate the identifying assumption as parallel trends across all vote shares, i.e., $E[Y^{0}_{j,k\geq 0}-Y^{0}_{j,k<0}|V_{j}]=E[Y^{0}_{j,k\geq 0}-Y^{0}_{j,k<0}]$, which yields the testable implication that trends should be parallel between losing elections with different vote shares. To implement this test, we modify the DiD model as follows:
@@
where $\nu^{g}$ denotes a complete set of vote share categories. In particular, we divide the vote share distribution into the following six groups: 0-20%, 20-35%, 35-50%, 50-65%, 65-80%, 80-100%. In the model we omit the 20-35% vote share category, such that all estimated effects must be interpreted relative to that group. Significant estimates for the 0-20% or 35-50% categories would then indicate the presence of unobserved shocks that drive both voting results and campaign contribution behavior.
@@
Third, we relax the parallel trends assumption by restricting the sample to elections where the union won or lost by an increasingly close margin. Establishments with closer election results can be expected to be more similar not only in terms of baseline characteristics but also in terms of shocks that they are exposed to over time. Specifically, we examine the robustness of the DiD estimates when restricting the sample to increasingly small vote share bandwidths around the 50% cutoff. In the limit, when comparing establishments where the union barely lost versus where it barely won, we approach the discontinuity-in-differences model estimated by Frandsen (2021) and Knepper (2020). For our baseline results from models (1) and (2), however, we follow Wang and Young (2023) and consider all elections with a pro-union vote share between 20% and 80%. This improves power and allows us to generalize effects for a broader sample of union elections.
Alternative source of variation. Lastly, we describe our DiD-IV approach, which complements the DiD strategy with arguably exogenous variation in union support driven by spikes in work-related fatalities. After the NLRB accepts a petition to hold a union election, it sets the timeline of the unionization process and fixes an election date. Any random unexpected shocks between petition and election that shift union support are then potential candidates for an instrument. We focus on sector-level fatal work accidents in the 30 days before a union election. [^27] The median time between petition and election in our sample is 45 days. Less than 5% of all elections are held within 30 days after the petition. Safety at work is a fundamental concern to all workers, especially when one’s life is in danger. Work-related fatalities are unfortunately still common in the United States. In 2019, the Occupational Safety and Health Administration (OSHA) reported 1,943 deaths at work, more than 5 per day on average. Unions often campaign on safety issues and are found to improve safety conditions in the workplace (e.g., AFL-CIO, 2022; Hagedorn et al., 2016; Li et al., 2022).
We implement the DiD-IV approach by estimating the following two-stage model:
where $\Delta y_{j}$ denotes the change in campaign contribution patterns in the three cycles after the union election relative to the three cycles before (excluding the cycle of the union election). By using changes as the outcome variable, the specification builds on the DiD approach and accounts for time-invariant differences between establishments that may affect the level of campaign contributions. Our main instrument is $A_{st}$, which represents the number of fatal accidents in 2-digit sector $s$ in the 30 days prior to the election after accounting for seasonal variation. [^28] Data on fatal work accidents is obtained from OSHA in the form of Fatality and Catastrophe Investigation Summaries (OSHA form 170). Appendix Figure A.3 depicts the exploited time variation. We allow for a non-linear effect by including $A_{st}^{2}$ and for a larger impact of fatalities in sectors where fatalities are common and where workers may be more concerned about workplace safety by the interaction term $A_{st}\times FR_{s}$ ($FR_{s}$ denotes the share of fatal work accidents occurring in a given sector out of all fatal work accidents in the sample). Importantly, instead of directly instrumenting union victory in a standard 2SLS approach, the first stage explains the continuous pro-union vote share $V_{j}$. In the second stage, we then use an indicator for predicted victory that is based on the predicted vote share in the first stage. This approach resembles the treatment assignment process and exploits the maximum available information. To account for the uncertainty from the first-stage regression, we compute standard errors by bootstrapping the two-stage model. In addition, we include several control variables. First, we account for the main effect of $FR_{s}$. Second, we include the yearly number of fatalities in a sector to ensure that sector-specific trends in fatalities do not drive our results. Third, we add the log number of employees at the sector-year level and the log number of eligible voters as precision controls. Finally, we include year fixed effects $\gamma_{t}$ and month-of-the-year fixed effects $\mu_{m}$.
The exclusion restriction of the instrument relies on the notion that a shock in fatal work accidents in the same sector affects political behavior only through its impact on the likelihood that an establishment will unionize. Two points are worth highlighting in that regard. First, all individuals in our sample are potentially exposed to the information on fatal work accidents. However, only some vote on unionization in the following 30 days. That is to say, we do not exploit differences in the direct exposure to work accidents but differences in the timing of the union election relative to the information shock. Second, we are solely focusing on the medium-term impact of spikes in fatal accidents by considering campaign contributions in the three cycles after the union election. The result that common shocks in fatal work accidents influence political behavior years afterward in some but not other establishments would be difficult to explain other than through the path dependency triggered by the increase in the likelihood of unionization shortly after the accidents.
@@
Definition of outcome variables. We consider two primary outcomes of employees’ political behavior at the establishment level. The first is the total amount of campaign contributions to all political candidates which we interpret as a measure of employees’ political participation and mobilization. Our second main outcome is the difference between the contribution amounts to Democratic and Republican candidates. Given the extant evidence on ideological motivations driving individuals’ donation behavior, we interpret the party composition of donations as a measure of employees’ ideological positions.
@@
In our baseline specification, we transform contribution amounts using the inverse hyperbolic sine (IHS) function to approximate log changes in amounts while retaining zero values. [^29] The inverse hyperbolic sine function is defined as $IHS(x)=ln(x+\sqrt{x^{2}+1})$. For sufficiently large $x$, $IHS(x)\approx ln(x)+ln(2)$. The function approximates the natural logarithm function for positive values but is well-defined for zero values. As demonstrated in recent econometric work (Aihounton and Henningsen, 2021; Chen and Roth, 2024; Mullahy and Norton, 2024), the IHS transformation yields marginal effects that can depend on the units in which the outcome is measured. To avoid arbitrary scaling of our outcomes, we always measure contribution amounts in 2010 USD before applying the IHS transformation. Chen and Roth (2024) show that the scale dependence arises because the scale implicitly determines the value that IHS assigns to changes along the extensive versus intensive margin. If the treatment has larger extensive margin effects – in our case, if unionization induces employees to donate positive amounts – the estimates using IHS are more sensitive to scaling. In section 5.3, we thus analyze the extensive margin effects of unionization and check the robustness when explicitly assigning alternative weights to extensive versus intensive margin changes.
@@ 5.1 Main Results
Figure 2 reports our main results for the effect of unionization on employees’ campaign contributions, using the DiD and event study models (1) and (2). We start with the effects on the (IHS-transformed) total contribution amounts, which are depicted on the left-hand side of the figure. The upper panel plots the results for all employees in an establishment. Note the absence of any significant differential trends between establishments winning and establishments losing the union election in the three cycles (six years) before the election. The effect of unionization on the amount of contributions is small and insignificant in all post-election periods, but we see a moderate spike in contributions in the cycle of the union election (which we are not able to estimate precisely, though). Differentiating between contributions made by workers and managers in the lower panels highlights that workers drive the increase in contributions. The event-study results indicate that unionization raises workers’ contributions by 11% in the cycle of the union election (significant at the 5% level). This pattern is consistent with a short-term political mobilization of workers through a successful union campaign in the workplace. Overall, however, the DiD coefficients indicate that there is no significant average effect on the amount of contributions over the three cycles after a union election.
@@
Next, we assess changes in the party composition of campaign contributions. If unions are able to change individuals’ political views, campaign contributions will shift to different candidates. The right-hand side of Figure 2 plots the effect of unionization on the difference between the (IHS-transformed) amounts spent to Democratic versus Republican candidates. First focusing on all employees, we again see no differential trends in contribution composition before the election. After the election, however, there is a significant increase in contributions donated to Democratic relative to Republican candidates. The effect on partisan support appears to be strongest in the long term, i.e., six years after the election. The DiD coefficient indicates that, over all post-election periods, unionization increases the difference in contributions to Democrats versus Republicans by 24 percentage points (significant at the 1% level). Differentiating again between workers and management in the lower two panels reveals that the effect is driven similarly by both groups. Not only workers but also managers significantly shift contributions from Republican to Democrat candidates in response to successful unionization. Quantitatively, the DiD estimates show that winning the union election increases donations to Democrats relative to Republicans by 12 percentage points for workers and by 20 percentage points for managers (both significant at the 1% level). These patterns are not consistent with an increase in tensions between unionized workers and their management but rather point toward an alignment of ideological positions.

### v5

@@ 4 Empirical Strategy
Our objective is to estimate the causal effect of unionization on employees’ campaign contributions. A simple comparison of employees in unionized and non-unionized workplaces will fail to account for differences between these groups along a number of dimensions. These arise because the decision to unionize is likely endogenous and correlated with many characteristics, among them potentially political behavior. Figure 1 depicts average contribution amounts and their party composition across won and lost union elections before and after the election. Due to their shared interest in a union election at the same time, these establishments are expected to be more similar than a random sample of unionized and non-unionized establishments. [^26] Dinlersoz2017 examine selection into union elections and find that elections are more likely to be held at younger, larger, more productive, and higher-paying establishments. Our strategy avoids such selection by comparing only establishments that hold union elections. Pre-existing ideological differences are nevertheless visible: Workplaces that vote for unionization donate more to Democratic relative to Republican candidates even before the union election.
@@
To account for pre-existing differences, we implement a difference-in-differences approach and compare campaign contribution patterns before and after the union election in establishments where the union won versus where it lost. We complement the DiD design with methods from the RDD literature to probe the validity of the underlying parallel trends assumption. In particular, we leverage the pro-union vote share, which discontinuously determines unionization at the 50% threshold, to estimate placebo tests for differential trends by vote shares among losing union elections as well as to examine the robustness of our DiD estimates when restricting the sample to establishments with increasingly close election results. [^27] Many papers on the effects of unionization follow an RDD by comparing establishments in which the union barely won versus where it barely lost (e.g., Campello2018; DiNardo2004; Ghaly2021; Lee2012; Sojourner2015a; Sojourner2015). This approach is complicated by the ability of both unions and employers to influence election outcomes even after the election, either by challenging the validity of individual ballots or by filing unfair labor practices charges. Frandsen2020 and Knepper2020 provide evidence for discontinuities at the 50% threshold in the vote share distribution and in pre-election establishment characteristics. Appendix Figure A.2 verifies that, also in our matched sample of elections, there is a significant discontinuity in the vote share density at the 50% cutoff, indicating a manipulation of closely contested elections. In addition, focusing on close elections may yield treatment effect estimates that are not representative for higher margin-of-victory elections. As Wang2021 show, barely won elections exhibit more decertification elections and delays in the election process, indicating a lower bargaining power of unions relative to elections with higher pro-union vote shares.
@@
where $y_{jk}$ denotes a political outcome for employees in establishment $j$ and relative event time $k$. We observe each establishment from three cycles before to three cycles after the union election, i.e., $k=\{-3,-2,...,3\}$, where $k=0$ refers to the cycle in which the union election takes place. Our effect of interest is captured by $\delta_{\textrm{DiD}}$. It is the coefficient of an interaction term between a post-treatment dummy and a dummy indicating whether the election was won by the union, i.e., whether the pro-union vote share, $V_{j}$, is above 50%. $\alpha_{j}$ denotes establishment fixed effects that capture all time-invariant differences between winning and losing establishments. Further, we introduce event-time $\times$ cohort fixed effects $\beta_{kc_{j}}$, where cohort $c_{j}$ refers to the political election cycle in which the union election was held, i.e., $c_{j}=\{1985/86,1987/88,...,2009/10\}$. Importantly, with these fixed effects our identifying variation only comes from comparing changes across won and lost elections within the same cohort. Thereby, our model is equivalent to the stacking approach used in staggered DiD settings (Cengiz2019) which avoids “forbidden comparisons” between late and early-treated establishments that may lead to negative weights when averaging potentially heterogeneous, cohort-specific treatment effects in these settings (DeChaisemartin2020; Goodman-Bacon2021; Sun2021). [^28] Note that in our case the selection of appropriate control units for each cohort of treated units in the stacking approach is facilitated by the possibility that we can naturally compare treated establishments with a won union election to untreated establishments that have a lost union election in the same cycle. Finally, we cluster standard errors at the establishment level, which is the unit of treatment assignment.
Model (1) pools all periods after treatment, which maximizes power when estimating average treatment effects. To examine how treatment effects vary by event time, we also estimate the following stacked event-study model:
@@
where the $\delta_{s}$ coefficients capture dynamic treatment effects relative to the cycle before the union election was held (the interaction with k=1k=-1 is omitted).
Parallel trends assumption. The identifying assumption is that campaign contributions for winning establishments would have evolved in parallel to contributions in losing establishments had the union not won the election:
$E[Y^{0}_{j,k\geq 0}-Y^{0}_{j,k<0}|V_{j}>.5]=E[Y^{0}_{j,k\geq 0}-Y^{0}_{j,k<0}|V_{j}\leq.5],$
@@
We run different tests to examine the validity of this assumption. First, we analyze whether outcomes developed in parallel before the election. Figure 1 provides first visual evidence that pre-election changes in total contribution amounts and in their party composition are very similar across won and lost elections. [^29] Note that the strong upward trends in total contribution amounts are explained by the increasing importance of campaign contributions in more recent election campaigns. The pre-election $\delta_{s}$ coefficients estimated in the event study model will provide a formal test of pre-trends.
Second, even in the absence of significant pre-trends, there may still be unobserved shocks that drive union voting results at the time of the election and that are related to changes in contribution patterns. To test whether such shocks likely violate our identifying assumption, we follow the approach of Wang2021 and analyze whether changes in outcomes are different among losing elections with different vote shares. If unobserved shocks were driving voting results that led to union victory or loss, we would also expect them to affect outcomes in losing elections with different union vote shares. [^30] Wang2021 formulate the identifying assumption as parallel trends across all vote shares, i.e., $E[Y^{0}_{j,k\geq 0}-Y^{0}_{j,k<0}|V_{j}]=E[Y^{0}_{j,k\geq 0}-Y^{0}_{j,k<0}]$, which yields the testable implication that trends should be parallel between losing elections with different vote shares. To implement this test, we modify the DiD model as follows:
@@
where $\nu^{g}$ denotes a complete set of vote share categories. In particular, we divide the vote share distribution into the following six groups: 0-20%, 20-35%, 35-50%, 50-65%, 65-80%, 80-100%. In the model we omit the 20-35% vote share category, such that all estimated effects are interpreted relative to that group. Significant estimates for the 0-20% or 35-50% categories would then indicate the presence of unobserved shocks that drive both voting results in the union election and campaign contribution behavior.
@@
Third, we restrict the sample to elections where the union won or lost by an increasingly close margin. Establishments with closer election results can be expected to be more similar not only in terms of baseline characteristics but also in terms of shocks that they are exposed to over time. Specifically, we examine the robustness of the DiD estimates when narrowing the sample to increasingly small vote share bandwidths around the 50% cutoff. In the limit, when comparing establishments where the union barely lost versus where it barely won, we approach the discontinuity-in-differences model estimated by Frandsen2020 and Knepper2020. For our baseline results from models (1) and (2), however, we follow Wang2021 and consider all elections with a pro-union vote share between 20% and 80%. This improves power and allows us to generalize effects for a broader sample of union elections.
@@
Definition of outcome variables. Throughout the analysis, we consider two primary outcomes of employees’ political behavior at the establishment level. The first is the total amount of campaign contributions to all political candidates which we interpret as a measure of employees’ political participation and mobilization. Our second main outcome is the difference between the contribution amounts to Democratic and Republican candidates. Given the extant evidence on ideological motivations driving individuals’ donation behavior, we interpret the party composition of donations as a measure of employees’ ideological positions.
@@
In our baseline specification, we transform contribution amounts using the inverse hyperbolic sine (IHS) function to approximate log changes in amounts while retaining zero values. As demonstrated in recent econometric work (Aihounton2021; Chen2024; Mullahy2024), the IHS transformation yields marginal effects that can depend on the units in which the outcome is measured. To avoid arbitrary scaling of our outcomes, we always measure contribution amounts in 2010 USD before applying the IHS transformation. Chen2024 show that the scale dependence arises because the scale implicitly determines the value that IHS assigns to changes along the extensive versus intensive margin. If the treatment has larger extensive margin effects – in our case, if unionization induces employees to donate positive amounts – the estimates using IHS are more sensitive to scaling. In section 5.3, we thus analyze the extensive margin effects of unionization and check the robustness when explicitly assigning alternative weights to extensive versus intensive margin changes. [^31] Following the recommendation of Chen2024, explicitly weighting extensive and intensive margin changes allows us to value a percentage point change in campaign contributions equally for all individuals, regardless of their initial contribution levels. Intuitively, a 10 USD increase in contributions by a lower-income individual may reflect the same ideological shift as a 100 USD increase by a more affluent individual. Alternative estimation strategies that avoid scale dependence by estimating treatment effects in levels fail to capture decreasing marginal utilities and may understate ideological shifts among lower-income donors.
@@ 5.1 Main Results
Figure 2 reports our main results for the effect of unionization on employees’ campaign contributions, using the DiD and event study models (1) and (2). We start with the effects on the (IHS-transformed) total contribution amounts, which are depicted in the left-hand panels of the figure. The upper panel plots the results for all employees in an establishment. Note the absence of any significant differential trends between establishments winning and establishments losing the union election in the three cycles (six years) before the election. The effect of unionization on the amount of contributions is small and insignificant in all post-election periods, but we see a moderate spike in contributions in the cycle of the union election (which we are not able to estimate precisely, though). Differentiating between contributions made by workers and managers in the lower panels highlights that workers drive the increase in contributions. The event-study results indicate that unionization raises workers’ contributions by 11% in the cycle of the union election (significant at the 5% level). This pattern is consistent with a short-term political mobilization of workers through a successful union campaign in the workplace. Overall, however, the DiD coefficients indicate that there is no significant average effect on the amount of contributions over the three cycles after a union election.
@@
Next, we assess changes in the party composition of campaign contributions. If unions are able to change individuals’ political views, campaign contributions will shift to different candidates. The right-hand panels of Figure 2 plot estimates for the effect of unionization on the difference in (IHS-transformed) amounts spent to Democratic versus Republican candidates. Starting with all employees, we again see no differential trends in contribution composition before the election. After the election, however, there is a significant increase in contributions donated to Democratic relative to Republican candidates. The partisan shift appears to be strongest in the long term, i.e., six years after the union election. The DiD estimate indicates that, over all post-election periods, unionization increases the Democratic-Republican contribution gap by 24 percentage points (significant at the 1% level). Differentiating again between workers and workers in the lower two panels reveals that the effect occurs among both groups. Not only workers but also managers significantly shift contributions from Republican to Democrat candidates in response to successful unionization. Quantitatively, the DiD estimates show that winning the union election increases donations to Democrats relative to Republicans by 12 percentage points for workers and by 20 percentage points for managers (both significant at the 1% level). [^32] Appendix Table A.4 also reports results separately for amounts contributed to Democratic and Republican candidates. For both workers and managers, the coefficients for contributions to Democrats are positive and those for Republicans are negative. However, among workers only the increase in contributions to Democrats is statistically significant, while among managers only the decrease in contributions to Republicans is significant. That said, our main results show that these party differences are not large enough to generate significant effects on total contribution amounts. The relatively stronger response among managers may seem surprising, as they are not the beneficiaries of unionization and are typically not targeted by unions’ political activities. One possible explanation is that managers begin from more conservative ideological positions – see, again, the average donation patterns in Table 2 – which increases the scope for changes after unionization. Workers, by contrast, already show strong Democratic support prior to unionization, leaving less room for change. Overall, our results point a political realignment within the workplace, where unionization narrows ideological gaps between managers and their unionizing workers.